OpenStudy (anonymous):

How many solutions will the following equation have? 25=20x-4x^2 A.0 B.1 C.2 D.3

OpenStudy (anonymous):

1 solution

thomaster (thomaster):

@TjRoDz Don't give the answer... guide people to the answer

OpenStudy (anonymous):

ok sorry

thomaster (thomaster):

\(\Large -4x^2+20x=25\) \(\Large -4x^2+20x-25=0\) a= -4 b=20 c=-25 \(\Large D=b^2-4ac\) \(\Large D=20^2-4*-4*-25=0\) When D (the discriminant) is zero, the equation has 1 solution

OpenStudy (anonymous):

ok cool